If the atomic weight of carbon to be 6 amu the value of Avogadro constant will be. Exp with steps.

Grade:12th pass

2 Answers

Harishwar IIT Roorkee
8 years ago
The number 6.02214179×1023alone is calledAvogadro's number (NA) orAvogadro'sconstant . It wont be changed with change in Carbon’s atomic mass
Tanmay
6 years ago
If at.wt. of carbon is 12u then AVG no.is 6.02×10^23Then 6u will give AVG no=6.02×10^23÷12×6=3.01×10^23